In a large vase, there are 8 roses, 5 daisies, 12 lilies, and 9 orchids. If 4 flowers are selected at random, find the probability that at least one of the flowers is a rose.
Thank you very much

Answer by stanbon(75887)

In a large vase, there are 8 roses, 5 daisies, 12 lilies, and 9 orchids. If 4 flowers are selected at random, find the probability that at least one of the flowers is a rose.
---
P(x>=1) = 1 - P(x=0)
= 1 - [26C6/34C4]
---
= 1 - 0.3224
---
= 0.6776
=================
Cheers,
Stan H.